OpenStudy (anonymous):

Solve for x. 4 • 7 – 4 • 10 = 4(x – 10)

OpenStudy (solomonzelman):

\[4 \times 7 -4 = 4(x-10)\]expand the right side, and you get \[4 \times 7 - 4 \times 10=4 \times x - 4 \times 10\]\[28-40=4x-40\]\[28=4x~~~~~~~~->~~~~~~~~x=7\] this was a standard approach now the way I would do it is. \[4 \times 7 - 4 \times 10= 4(x-10)\]factor the left side,\[4(7-10)=4(x-10)\]\[4(\color{blue} { 7 } -10)=4(\color{blue} { x } -10)\] see what the x is equal to? (just a little shortcut)